OpenStudy (shiree19):

I have a two questions. How do you do these types of problems and what are their answers? *Is -7 + 9 = -9 + 7 true, false, or open? **Is 4x - 3 = 19 true, false, or open? Thanks.

OpenStudy (math&ing001):

You do these types of problems by calculating : * -7 + 9 = -9 + 7 => 2=-2 So it's false. ** 4x - 3 = 19 => x=22/4 So this is open (only true when x=11/2 )

OpenStudy (shiree19):

oh ok. so I do each side and if they are equal it's true. If not it's false. If it has a letter (algebraic equation) it's open.

OpenStudy (math&ing001):

Yep, and you solve for that letter first, if there is a solution then it's open if there isn't you check if it's true or false.

OpenStudy (shiree19):

ok. thanks so much!

OpenStudy (math&ing001):

Welcome !